# Data Science for Geographic Information Systems

### v1

@@ 2 Keywords:
3 Introduction
@@
This work is divided into six sections as follows: Firstly, a historical record of the development of the fields of data science (Section 4) and of GISs (Section 5). Then, an overview of the convergence of the two fields (Section 6) followed by an illustration of the various applications in different areas of interest that hinge on this convergence (Section 7). Finally, we proceed with a small case study to illustrate the process of extracting insights from raw data (Section 8), and conclude with some final remarks on the work and prospects on the importance of this synergy for future work (Section 9).
4 Emergence of Data Science
@@
The evolution of data science can be traced through transformative milestones. In its early stages, as machine learning gained traction (Flach, 2012; Borges, 1989) and more sophisticated statistical methods were developed (Kotz and Johnson, 2012; Freedman, 2009), relational databases and data warehousing technologies became crucial for managing and organizing data (Chaudhuri and Dayal, 1997).
@@
With the concept of data warehousing gaining prominence, emphasizing the need for centralized repositories to store integrated historical data (Gardner, 1998; Kimball, 1996), advances in data mining (Fayyad et al., 1996) leveraged these emerging technologies as the backbone for storing and managing larger datasets (Chaudhuri and Dayal, 1997).
@@
Eventually, the emergence of big data (Mayer-Schönberger and Cukier, 2013), accompanied by the widespread adoption of open-source tool ecosystems, such as R and Python (Ihaka and Gentleman, 1996; McKinney, 2012), reached a critical point. Later data science would evolve into a recognized, specialized, and interdisciplinary field (Dhar, 2013), witnessing the big resurgence of deep learning and its integration across industries (Goodfellow et al., 2016), significantly impacting business decision-making (Schoenherr and Speier-Pero, 2015; Van Der Aalst and van der Aalst, 2016).
@@
Recently, there has been an increasing focus on ethical considerations, responsible AI practices, and fairness in data science applications (Diakopoulos, 2016; Mittelstadt et al., 2016). Regardless, data science remains a vital component of decision-making processes in most industries, reflecting the dynamic nature of its continuous development over time.
5 Evolution of Geographic Information Systems
@@
A GIS is a system designed to capture, store, analyze, manage, and present spatial data. GIS allows users to visualize, interpret, and understand data in a spatial context (Longley et al., 2015; Burrough et al., 2015). Since their inception in the early 1960s, these systems have undergone a notable evolution. Originally conceived simply as computerized tools for land resource management, GIS has progressed through various stages of development.
@@
During its initial stages, GIS primarily concentrated on the digitization of maps and the conceptualization of spatial databases. The foundational work carried out by Roger Tomlinson (Tomlinson, 1974) set the stage for the development of the first computerized GIS. With advances in computer technology, the 1970s and early 1980s witnessed the commercialization of GIS software, with significant contributions from companies such as the Environmental Systems Research Institute (ESRI).
@@
During the following decades, the incorporation of remote sensing data—satellite imagery in particular—was a pivotal development for enhancing the capabilities of GIS (Merchant, 2000). Standardization efforts, led by the Open Geospatial Consortium (OGC), aimed to establish interoperability among GIS platforms. Furthermore, the 1990s also included the first definitions of GIS as a scientific area of development—GIScience (Rocha and Abrantes, 2019).
@@
During the early 2000s, GIS underwent a transition to the web, facilitating greater accessibility to spatial data (Chang, 2008). Subsequently, the rise of cloud-based GIS platforms and the popularity of web mapping applications (Peterson, 2005), such as Google Maps, significantly contributed to the widespread recognition of GIS.
@@
The 2010s and beyond witnessed the advent of mobile GIS and the growth of location-based services (Hussein et al., 2011), harnessing the widespread use of smartphones for real-time data collection and field mapping (Khan et al., 2012). During this time, there was also a considerable adoption of big data methods and advanced spatial analytics within GIS applications (Goodchild, 2016).
6 Key Components of Data Science in GIS
@@
This section explores and covers key aspects of the synergy between GIS and data science. It begins with spatial data handling and addresses techniques and challenges in preprocessing and analysis ( Subsection 6.1). Next, the impact of Machine Learning (ML) in spatial analysis is explored ( Subsection 6.2), followed by several insights on the use of Big Data technologies in the context of GIS and the importance of scalable solutions ( Subsection 6.3). Finally, the section concludes with the importance of effective data communication in GIS, offering a brief explanation on spatial data representation and interpretation ( Subsection 6.4).
6.1 Spatial Data Handling
@@
Spatial data handling in GIS involves the application of data science techniques to efficiently manage and process geographical information. Techniques, including spatial indexing, geocoding, and spatial querying, play important roles in optimizing data retrieval and analysis. Spatial Indexing involves organizing data to facilitate efficient spatial queries, optimizing data retrieval processes. Geocoding converts places of interest into geographic coordinates, a fundamental step for mapping and essential for spatial analysis and data integration. Lastly, Spatial Querying allows for the retrieval of information based on spatial relations or conditions, offering valuable insights from spatial data crucial for decision-making processes. This field is deemed so crucial that it is often considered a subset of data science itself—spatial data science (Garrard, 2016; Rey et al., 2023; Comber and Brunsdon, 2020).
@@
6.2 Machine Learning
ML has emerged as a game-changing tool within the domain of GIS, facilitating the extraction of valuable insights from spatial datasets and augmenting traditional spatial analysis (Choudhury and Um, 2023; Casali et al., 2022). ML algorithms brought a data-driven and adaptive approach to spatial problem-solving, allowing GIS professionals to uncover more intricate patterns, make more accurate predictions, and automate once complex and time consuming tasks.
Feature recognition and extraction are ML techniques that allow automating the identification of spatial patterns and objects. This capability is particularly evident in applications like image classification (Sulemane et al., 2022; Lim et al., 2019), where ML algorithms discern and categorize features within satellite or aerial imagery, contributing to tasks such as land-use classification and environmental monitoring.
6.3 Big Data
@@
To address the challenges presented by large spatial datasets, GIS developers are increasingly embracing Big Data technologies and methodologies (Akhund, 2022; Lee and Kang, 2015; Li et al., 2016). This shift aims to establish scalable and distributed computing frameworks that efficiently process extensive amounts of spatial data. Two critical paradigms in this context are the parallelization of spatial data (Zhao et al., 2016; Werner, 2019) and the scalability of spatial systems (Jhummarwala et al., 2014; Neuman, 1994; Hawick et al., 2003).
@@
The significance of scalable and efficient data processing in the context of Big Data and GIS cannot be overstated. As spatial datasets continue to grow, the ability to process information at scale becomes paramount for timely decision-making and insightful analysis. Scalable data processing not only addresses the challenges of handling large datasets more efficiently but also unlocks the potential for real-time (Sun and Li, 2016; Al-Sabhan et al., 2003) and near-real-time (Gamba and Casciati, 1998) spatial analytics, enabling a dynamic response to evolving spatial phenomena previously difficult to achieve (Zerger and Smith, 2003).
6.4 Data Visualization and Interpretation
Effective communication of spatial information is a key aspect of GIS, and data visualization is a powerful tool in achieving this goal (Longley et al., 2015). Data science and data visualization techniques are employed to create visually appealing and informative representations, allowing the transformation of large and complex spatial datasets into comprehensible and insightful depictions (Ki, 2018).
Well-designed maps and graphics enhance the interpretability of spatial information (Slocum et al., 2022), enabling stakeholders to make more informed decisions. GIS enables the creation of a plethora of spatial data representations—whether in 2D, 3D, or even VR—facilitating the analysis and conveyance of complex spatial patterns and trends in an accessible manner (Mitchell, 1999; Mitchel et al., 2005).
@@ 7 Applications
The integration of GIS and data science has spurred transformative applications across various domains. This section will explore some key areas where the synergy of GIS and data science has been notable, highlighting the versatility and impact of these fields. We start by looking into urban and infrastructure planning (Subsection 7.1) and examining the far-reaching implications on environmental analysis and conservation efforts (Subsection 7.2). Moving forward, we explore the impact on public health and safety (Subsection 7.3). Finally, we discuss several applications within the agricultural domain (Subsection 7.4). This exploration will underscore the multifaceted contributions of GIS and data science in addressing complex challenges across various sectors.
7.1 Urban and Infrastructure Planning
@@
By analyzing large datasets, urban planners can gain valuable insights into population trends, traffic patterns (Yang et al., 2019), and public infrastructure usage (Mohammadi and Al-Fuqaha, 2018). This information aids in designing efficient transportation systems (Torre-Bastida et al., 2018) and optimizing current city layouts and future developments (Hao et al., 2015). Furthermore, machine learning algorithms can be employed to predict urban growth, enabling proactive planning to accommodate increasing population demands (Chaturvedi and de Vries, 2021). Additionally, GIS-based data science has the potential to help in assessing the impact of urban projects on the environment and social dynamics (Kitchin, 2016).
7.2 Environmental Analysis and Conservation
@@
Remote sensing data, satellite imagery, and ground-based sensor data can be integrated to monitor dynamic changes in ecosystems (Hansen and Loveland, 2012), enabling the identification of deforestation and unexpected alterations in habitats (Pettorelli et al., 2014; Sexton et al., 2013). Furthermore, advanced analytics facilitate the discernment of environmental patterns, allowing conservationists to make well-informed decisions regarding conservation efforts (Horning et al., 2010). Finally, using historical data, machine learning models can help predict the occurrence and impact of natural disasters (Linardos et al., 2022), including wildfires (Jain et al., 2020; Oliveira et al., 2023; Fachada, 2022), earthquakes (Galkina and Grafeeva, 2019), or floods (Mosavi et al., 2018). This predictive capability enhances preparedness and response strategies for more effective disaster management.
7.3 Public Health and Safety
@@
In the domain of public health, integrated demographic information, disease prevalence, and environmental factors data can be used to identify potential health risks and patterns (Kubben et al., 2019). Maps can be employed to visualize the spread of diseases, plan healthcare infrastructure, and optimize resource allocation during health emergencies (Kamel Boulos and Geraghty, 2020). Predictive modeling aids in forecasting disease outbreaks, enabling authorities to implement efficient interventions (Pullan et al., 2011). Additionally, spatial data science can contribute to identify environmental factors affecting public health, supporting initiatives for cleaner and healthier living environments (Maantay and McLafferty, 2011).
7.4 Agricultural Optimization
@@
Data science applications in GIS can significantly enhance agricultural practices by providing insights into crop health, soil conditions, and weather patterns (Wardlow et al., 2007; Mitran et al., 2021). Monitoring crop growth, detecting diseases, and optimizing irrigation become feasible through the integration of satellite imagery and terrain sensor data, coupled with predictive algorithms (Thenkabail et al., 2009). Furthermore, these applications can predict crop yields, identify optimal planting times, determine suitable crop types, and recommend maintenance methods (He et al., 2017; Priya and Shibasaki, 2001; dos Santos et al., 2023).
Figure 2: Case study results, with a scale of 1 pixel per square meter: a) RGB representation of area of study, in Tróia, Portugal; b) slope in degrees; c) aspect (orientation of slope) in degrees; d) elevation in meters above sea level; e) canopy height in meters; f) vegetation enhanced; g) water features enhanced, with attention drawn to a subtle disturbance at the lower left corner, near the coastline; h) man-made structures; i) fuel map with fuel codes corresponding to a model developed based on (Fernandes and Loureiro, 2022); and, j) burning Index map generated through simulations using an additional tool and historical weather data.
8 Case Study
@@
For the purpose of this case study, we use two datasets: 1) aerial imagery encompassing red, blue, green, near-infrared (NIR), and red-edge bands; and, 2) a Digital Elevation Model (DEM). The focus of this study is a small land patch situated in Tróia, Portugal, an RGB representation of which is shown in Figure 2a. All data was collected in 2022.
@@
Finally, synthesizing all this knowledge, we can create a fuel map of the area, which, in turn, can serve diverse purposes, such as wildfire prevention measures through simulated prediction (Fachada, 2022). This could, for example, allow to the computation of a Burning Index (BI) map, serving as a danger indicator to assist authorities in making better decisions regarding wildfire combat and terrain maintenance. Figures 2i and 2j highlight the fuel and BI result for the area analyzed in this case study.
9 Conclusions
@@
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
Afonso Oliveira: Investigation, Methodology, Software, Writing - Original Draft. Nuno Fachada: Supervision, Writing - Review & Editing, Visualization. João P. Matos-Carvalho: Supervision, Writing - Review & Editing.
Funding
This work was financed by the Portuguese Agency FCT (Fundação para a Ciência e Tecnologia), in the framework of projects UIDB/00066/2020, UIDB/04111/2020 and CEECINST/00147/2018/CP1498/CT0015.
Data Availability Statement
Datasets are available on request. The raw data supporting the conclusions of this article will be made available by the authors, without undue reservation.

### v3

@@ Abstract
{textblock*}
190mm(-1cm,-6cm) The peer-reviewed version of this paper is published in IEEE Xplore at https://doi.org/10.1109/YEF-ECE62614.2024.10624902. This version is typeset by the authors and differs only in pagination and typographical detail.
I Introduction
@@
This work is divided into six sections as follows: Firstly, a historical record of the development of the fields of data science (Section II) and of GISs (Section III). Then, an overview of the convergence of the two fields (Section IV) followed by an illustration of the various applications in different areas of interest that hinge on this convergence (Section V). Finally, we proceed with a small case study to illustrate the process of extracting insights from raw data (Section VI), and conclude with some final remarks on the work and prospects on the importance of this synergy for future work (Section VII).
II Emergence of Data Science
@@
The evolution of data science can be traced through transformative milestones. In its early stages, as machine learning gained traction [1, 2] and more sophisticated statistical methods were developed [3, 4], relational databases and data warehousing technologies became crucial for managing and organizing data [5].
@@
With the concept of data warehousing gaining prominence, emphasizing the need for centralized repositories to store integrated historical data [6, 7], advances in data mining [8] leveraged these emerging technologies as the backbone for storing and managing larger datasets [5].
@@
Eventually, the emergence of big data [9], accompanied by the widespread adoption of open-source tool ecosystems, such as R and Python [10, 11], reached a critical point. Later data science would evolve into a recognized, specialized, and interdisciplinary field [12], witnessing the big resurgence of deep learning and its integration across industries [13], significantly impacting business decision-making [14, 15].
@@
Recently, there has been an increasing focus on ethical considerations, responsible AI practices, and fairness in data science applications [16, 17]. Regardless, data science remains a vital component of decision-making processes in most industries, reflecting the dynamic nature of its continuous development over time.
III Evolution of Geographic Information Systems
@@
A GIS is a system designed to capture, store, analyze, manage, and present spatial data. GIS allows users to visualize, interpret, and understand data in a spatial context [18, 19]. Since their inception in the early 1960s, these systems have undergone a notable evolution. Originally conceived simply as computerized tools for land resource management, GIS has progressed through various stages of development.
@@
During its initial stages, GIS primarily concentrated on the digitization of maps and the conceptualization of spatial databases. The foundational work carried out by Roger Tomlinson [20] set the stage for the development of the first computerized GIS. With advances in computer technology, the 1970s and early 1980s witnessed the commercialization of GIS software, with significant contributions from companies such as the Environmental Systems Research Institute (ESRI).
@@
During the following decades, the incorporation of remote sensing data—satellite imagery in particular—was a pivotal development for enhancing the capabilities of GIS [21]. Standardization efforts, led by the Open Geospatial Consortium (OGC), aimed to establish interoperability among GIS platforms. Furthermore, the 1990s also included the first definitions of GIS as a scientific area of development—GIScience [22].
@@
During the early 2000s, GIS underwent a transition to the web, facilitating greater accessibility to spatial data [23]. Subsequently, the rise of cloud-based GIS platforms and the popularity of web mapping applications [24], such as Google Maps, significantly contributed to the widespread recognition of GIS.
@@
The 2010s and beyond witnessed the advent of mobile GIS and the growth of location-based services [25], harnessing the widespread use of smartphones for real-time data collection and field mapping [26]. During this time, there was also a considerable adoption of big data methods and advanced spatial analytics within GIS applications [27].
IV Key Components of Data Science in GIS
@@
This section explores and covers key aspects of the synergy between GIS and data science. It begins with spatial data handling and addresses techniques and challenges in preprocessing and analysis ( Subsection IV-A). Next, the impact of Machine Learning (ML) in spatial analysis is explored ( Subsection IV-B), followed by several insights on the use of Big Data technologies in the context of GIS and the importance of scalable solutions ( Subsection IV-C). Finally, the section concludes with the importance of effective data communication in GIS, offering a brief explanation on spatial data representation and interpretation ( Subsection IV-D).
IV-A Spatial Data Handling
@@
Spatial data handling in GIS involves the application of data science techniques to efficiently manage and process geographical information. Techniques, including spatial indexing, geocoding, and spatial querying, play important roles in optimizing data retrieval and analysis. Spatial Indexing involves organizing data to facilitate efficient spatial queries, optimizing data retrieval processes. Geocoding converts places of interest into geographic coordinates, a fundamental step for mapping and essential for spatial analysis and data integration. Lastly, Spatial Querying allows for the retrieval of information based on spatial relations or conditions, offering valuable insights from spatial data crucial for decision-making processes. This field is deemed so crucial that it is often considered a subset of data science itself—spatial data science [28, 29, 30].
@@
IV-B Machine Learning
ML has emerged as a game-changing tool within the domain of GIS, facilitating the extraction of valuable insights from spatial datasets and augmenting traditional spatial analysis [31, 32]. ML algorithms brought a data-driven and adaptive approach to spatial problem-solving, allowing GIS professionals to uncover more intricate patterns, make more accurate predictions, and automate once complex and time consuming tasks.
Feature recognition and extraction are ML techniques that allow automating the identification of spatial patterns and objects. This capability is particularly evident in applications like image classification [33, 34], where ML algorithms discern and categorize features within satellite or aerial imagery, contributing to tasks such as land-use classification and environmental monitoring.
IV-C Big Data
@@
To address the challenges presented by large spatial datasets, GIS developers are increasingly embracing Big Data technologies and methodologies [35, 36, 37]. This shift aims to establish scalable and distributed computing frameworks that efficiently process extensive amounts of spatial data. Two critical paradigms in this context are the parallelization of spatial data [38, 39] and the scalability of spatial systems [40, 41, 42].
@@
The significance of scalable and efficient data processing in the context of Big Data and GIS cannot be overstated. As spatial datasets continue to grow, the ability to process information at scale becomes paramount for timely decision-making and insightful analysis. Scalable data processing not only addresses the challenges of handling large datasets more efficiently but also unlocks the potential for real-time [43, 44] and near-real-time [45] spatial analytics, enabling a dynamic response to evolving spatial phenomena previously difficult to achieve [46].
IV-D Data Visualization and Interpretation
Effective communication of spatial information is a key aspect of GIS, and data visualization is a powerful tool in achieving this goal [18]. Data science and data visualization techniques are employed to create visually appealing and informative representations, allowing the transformation of large and complex spatial datasets into comprehensible and insightful depictions [47].
Well-designed maps and graphics enhance the interpretability of spatial information [48], enabling stakeholders to make more informed decisions. GIS enables the creation of a plethora of spatial data representations—whether in 2D, 3D, or even VR—facilitating the analysis and conveyance of complex spatial patterns and trends in an accessible manner [49, 50].
@@ IV-D Data Visualization and Interpretation
(a) Choropleth Map, adapted from [51].
(b) Heat Map, adapted from [52]
(c) Flow Map, adapted from [53].
(d) 3D Height Model, adapted from [54]
@@ V Applications
The integration of GIS and data science has spurred transformative applications across various domains. This section will explore some key areas where the synergy of GIS and data science has been notable, highlighting the versatility and impact of these fields. We start by looking into urban and infrastructure planning (Subsection V-A) and examining the far-reaching implications on environmental analysis and conservation efforts (Subsection V-B). Moving forward, we explore the impact on public health and safety (Subsection V-C). Finally, we discuss several applications within the agricultural domain (Subsection V-D). This exploration will underscore the multifaceted contributions of GIS and data science in addressing complex challenges across various sectors.
V-A Urban and Infrastructure Planning
@@
By analyzing large datasets, urban planners can gain valuable insights into population trends, traffic patterns [55], and public infrastructure usage [56]. This information aids in designing efficient transportation systems [57] and optimizing current city layouts and future developments [58]. Furthermore, machine learning algorithms can be employed to predict urban growth, enabling proactive planning to accommodate increasing population demands [59]. Additionally, GIS-based data science has the potential to help in assessing the impact of urban projects on the environment and social dynamics [60].
V-B Environmental Analysis and Conservation
@@
Remote sensing data, satellite imagery, and ground-based sensor data can be integrated to monitor dynamic changes in ecosystems [61], enabling the identification of deforestation and unexpected alterations in habitats [62, 63]. Furthermore, advanced analytics facilitate the discernment of environmental patterns, allowing conservationists to make well-informed decisions regarding conservation efforts [64]. Finally, using historical data, machine learning models can help predict the occurrence and impact of natural disasters [65], including wildfires [66, 67, 68], earthquakes [69], or floods [70]. This predictive capability enhances preparedness and response strategies for more effective disaster management.
V-C Public Health and Safety
@@
In the domain of public health, integrated demographic information, disease prevalence, and environmental factors data can be used to identify potential health risks and patterns [71]. Maps can be employed to visualize the spread of diseases, plan healthcare infrastructure, and optimize resource allocation during health emergencies [72]. Predictive modeling aids in forecasting disease outbreaks, enabling authorities to implement efficient interventions [73]. Additionally, spatial data science can contribute to identify environmental factors affecting public health, supporting initiatives for cleaner and healthier living environments [74].
V-D Agricultural Optimization
@@
Data science applications in GIS can significantly enhance agricultural practices by providing insights into crop health, soil conditions, and weather patterns [75, 76, 77]. Monitoring crop growth, detecting diseases, and optimizing irrigation become feasible through the integration of satellite imagery and terrain sensor data, coupled with predictive algorithms [78, 79, 80]. Furthermore, these applications can predict crop yields, identify optimal planting times, determine suitable crop types, and recommend maintenance methods [81, 82, 83].
VI Case Study
@@
For the purpose of this case study, we use two datasets: 1) aerial imagery encompassing red, blue, green, near-infrared (NIR), and red-edge bands; and, 2) a Digital Elevation Model (DEM). The focus of this study is a small land patch situated in Tróia, Portugal, an RGB representation of which is shown in Figure 2a. All data was collected in 2022 [84, 85].
@@
Finally, synthesizing all this knowledge, we can create a fuel map of the area, which, in turn, can serve diverse purposes, such as wildfire prevention measures through simulated prediction [68]. This could, for example, allow to the computation of a Burning Index (BI) map, serving as a danger indicator to assist authorities in making better decisions regarding wildfire combat and terrain maintenance. Figures 2i and 2j highlight the fuel and BI result for the area analyzed in this case study.
Figure 2: Case study results, with a scale of 1 pixel per square meter: a) RGB representation of area of study, in Tróia, Portugal; b) slope in degrees; c) aspect (orientation of slope) in degrees; d) elevation in meters above sea level; e) canopy height in meters; f) vegetation enhanced; g) water features enhanced, with attention drawn to a subtle disturbance at the lower left corner, near the coastline; h) man-made structures; i) fuel map with fuel codes corresponding to a model developed based on [86]; and, j) burning Index map generated through simulations using an additional tool and historical weather data. The multispectral data was obtained from [87, 88].
VII Conclusions